# SAM2-ELNet: Label Enhancement and Automatic Annotation for Remote Sensing Segmentation

### v1

@@ Abstract
Remote sensing image segmentation is crucial for environmental monitoring, disaster assessment, and resource management, directly affecting the accuracy and efficiency of surface information extraction. The performance of existing supervised models in remote sensing image segmentation tasks highly depends on the quality of label data. However, current label data mainly relies on manual annotation, which comes with high time costs and is subject to subjective interference, resulting in distortion of label boundaries and often a loss of detail. To solve the above problems, our work proposes an Edge-enhanced Labeling Network, called SAM2-ELNet, which incorporates a labeling module and an edge attention mechanism. This model effectively addresses issues such as label detail loss, fragmentation, and inaccurate boundaries. Due to the scarcity of manually annotated remote sensing data, the feature extraction capabilities of traditional neural networks are limited. Our method uses the Hiera backbone of the pre-trained self-supervised large model segment anything model 2 (SAM2) as the encoder, achieves high-quality and efficient feature extraction even with small samples by fine-tuning on downstream tasks. This study compared the training effects of original and enhanced labels on the manually annotated Deep-SAR Oil Spill (SOS) dataset. Results showed that the model trained with enhanced labels performed better and had a lower final loss, indicating closer alignment with the real data distribution. Our work also explores the potential of extending the model into an efficient automatic annotation framework through generalization experiments, facilitating large-scale remote sensing image interpretation and intelligent recognition.
Index Terms:
Remote Sensing Image Segmentation, Automatic Annotation, Edge Attention Mechanism, Binarization Processing, Label Enhancement.
@@ I Introduction
Remote sensing typically refers to the technology of imaging the earth system in a specific electromagnetic spectrum band on an aerospace or aviation platform, so as to obtain multi-faceted characteristic information of the observed object. It’s imaging methods [1] include optical imaging, thermal infrared [2], hyperspectral [3], [4], and synthetic aperture radar (SAR) [5]. Currently, there are already various ways to extract valuable information from the massive amount of data. Remote sensing image segmentation has become a focal point of contemporary research due to its capability to delineate target regions and reveal terrestrial spatial structures.
Remote sensing image segmentation methods are mainly divided into traditional methods and deep learning-based methods. Traditional segmentation methods typically employ techniques such as thresholding, random forest [37], support vector machine [38], [16], and conditional random field [17] to delineate different surface structure regions. When applied to simple scenes, these methods often yield satisfactory results. But their sensitivity to radiometric noise, occlusive shadows, and topological ambiguities leads to issues, especially when dealing with high-dimensional features and complex spatial relationships.
Driven by advancements in the field of computer vision (CV), deep learning techniques have become the leading approach for remote sensing image segmentation [49], [50], [51], [52], [53]. Supervised learning methods in this domain rely on large-scale, high-quality annotated datasets to effectively train robust models. However, most existing datasets are manually annotated. The inherent heterogeneity of targets in remote sensing images, coupled with noise interference, affects boundary clarity, and increases the difficulty of precise recognition. Due to the subjectivity of annotators and the inherent complexity of the task, annotators often generate annotations of suboptimal quality during the labeling process, which tend to omit details or over-simplify geometric shapes[6]. As dataset scales expand exponentially, the corresponding annotation costs grow prohibitively. Therefore, reducing reliance on high-quality manual annotations, improving annotation efficiency and accuracy, and developing effective automated annotation frameworks are crucial for advancing intelligent remote sensing recognition and interpretation.
@@
With the growing adoption of these methods, researchers are seeking better ways to optimize manually annotated data while improving labeling efficiency in remote sensing applications. Compared with CV tasks, remote sensing images have more severe noise and scale changes, which makes end-to-end self-supervised methods face greater challenges. Wang et al. [60] proposed an automatic annotation tool that integrates multiple deep learning models (including SAM), called X-Anylabeling. Its application to remote sensing image annotation still requires manual input, and its effectiveness remains limited in scenarios involving small object detection, high-noise images, or abnormal data distributions. To enhance the application of SAM in remote sensing image segmentation, Zhang et al. [36] introduced the adapter-scale and adapter-feature modules to improve SAM’s encoding capability, making it more suitable for semantic segmentation tasks in remote sensing. Ma et al. [35] proposed the SGO-SGB framework by incorporating object consistency loss and boundary preservation loss, addressing SAM’s fragmentation and boundary inaccuracy issues in semantic segmentation of remote sensing imagery. Chen et al. [61] developed the RSPrompter method, which learns to generate category-specific prompts to enhance SAM’s applicability in remote sensing instance segmentation. Despite SAM2 performs well in many tasks, it may still generate segmentation results unrelated to specific categories when manual prompts are absent, which limits its usability in the specific context of remote sensing segmentation.
To address this limitation, Xiong et al. [15] integrated SAM2’s hierarchical feature extraction with U-Net’s spatial attention mechanism and proposed a novel network, SAM2-UNet. The architecture combines U-Net’s symmetrical decoder with a lightweight parameter adapter [62], [63], enabling efficient fine-tuning for downstream applications. Extending the SAM2-UNet framework, we propose a new model for label refinement and automatic annotation, called edge-enhanced labeling network (SAM2-ELNet). This model leverages large-scale self-supervised pretrained models for feature extraction and integrates few-shot learning for fine-tuning on downstream tasks. Our work introduce an edge attention mechanism and a labeling module to address issues such as label detail loss, fragmentation, and inaccurate boundaries in manually annotated remote sensing segmentation data, with the potential to be extended into an efficient automatic labeling framework.The main contributions of this study are as follows.
@@
    To address the issue of suboptimal annotations caused by heterogeneity and noise interference in remote sensing segmentation tasks, our work propose SAM2-ELNet, which incorporates a labeling module and an edge attention mechanism. This model effectively optimizes issues such as label detail loss, fragmentation, and inaccurate boundaries.
@@
    To enhance the applicability and training efficiency of SAM for specific remote sensing image segmentation tasks, we leverage a pre-trained large model and adopt a fine-tuning strategy with a frozen encoder. This approach retains the robust feature extraction capability of SAM2 while significantly improving training efficiency.
@@
    In exploring large-scale automatic annotation, we discuss the model’s generalization ability and investigate its potential to be extended into an efficient automatic labeling framework, facilitating the construction of scalable, high-quality datasets.
The remainder of this paper is organized as follows. Section II reviews related works on remote sensing image segmentation, automatic annotation methods, and the application of pretrained models in remote sensing. Section III presents the proposed SAM2-ELNet architecture, providing a detailed description of its components. Section IV offers a comprehensive experimental analysis, including model performance evaluation, ablation studies, generalization capability assessment, and efficiency comparisons. Both qualitative and quantitative analyses are conducted to validate the effectiveness of the proposed model. Finally, Section V concludes the study.
@@ II-A Advances in Annotation Strategies for Remote Sensing Image Segmentation
Yang et al.[18] introduced the EasySeg framework, which addresses domain adaptation challenges in remote sensing semantic segmentation. Their method specifically targets domain discrepancies and misperception issues in cross-domain applications. It incorporates a ’See-First-Ask-Later’ point-level annotation strategy and an Interactive Semantic Segmentation Network (ISS-Net), substantially reducing annotation costs while maintaining segmentation accuracy under weak supervision. Liu et al.[19] examined the role of noisy labels in supervised pretraining, analyzing their impact on encoder performance in remote sensing segmentation tasks. By comparing noisy label pretraining with self-supervised learning methods such as DINO and MoCo, they further demonstrated the effectiveness of noisy labels in pretraining and examined how class inconsistency influences model transferability. Li et al.[20] developed the SAM-OIL framework, which integrates YOLOv8-based object detection, an adapted SAM segmentation module, and an ordered mask fusion (OMF) mechanism for oil spill detection in SAR imagery. This method significantly improves detection accuracy while reducing dependence on fine-grained annotations.
@@ II-B Enhancing Annotation Quality and Efficiency with Advanced Learning Models
To further enhance annotation quality and efficiency, Wu et al.[21] pioneered the use of vision-language models (VLM) in farmland segmentation. They constructed a farmland image-text paired dataset and integrated semantic segmentation models with large language models (LLMs), developing the FSVLM model, which improved generalization and segmentation accuracy of remote sensing segmentation . Li et al.[22] introduced a semi-automatic annotation method that combines anchor real-time lines and ROI real-time line algorithms to address label incompleteness and deficiencies in traditional remote sensing indices and interactive segmentation. Han et al.[23] designed the DM-ProST framework, which generates high-quality pseudo-labels and expands training datasets through mutual correction between two deep learning models. This approach selects reliable unlabeled samples, improving segmentation accuracy. Zhu et al.[24] proposed AIO2 (adaptively triggered online object label correction), which integrates adaptive correction triggering (ACT) and online object label correction (O2C) to enhance robustness against noisy labels. Shi et al.[25] developed an automated multi-temporal change detection method for remote sensing images, incorporating boundary preservation techniques. By applying adaptive cropping and optimized classifier training, their method generates high-quality annotations and updates existing polygon-based labels. However, it does not resolve label omissions and struggles with non-polygonal targets such as oil spills and lakes.
@@ II-C Optimization and Application of SAM in Remote Sensing and Geoscience
The segment anything model has gained widespread attention for its role in self-supervised foundation models across remote sensing, medical imaging, and geophysical exploration[26]. In land use and crop type segmentation, SAM significantly improves performance[27], [28]. Ma et al.[29] explored geophysical foundation models (GeoFM) for geophysical exploration and applied SAM to seismic phase picking. SAM also enhances building segmentation and urban area delineation in remote sensing imagery [30], [31], [32], [33], [34], demonstrating its adaptability in processing complex geographical features.
Figure 1: Architecture of SAM2-ELNet. The proposed SAM2-ELNet comprises an encoder, decoder, receptive field blocks (RFB), and a labeling module. It integrates an edge attention mechanism and parameter-efficient fine-tuning (PEFT) to enhance segmentation accuracy and improve label quality for remote sensing imagery.
@@ III Proposed Method
Building upon the research advancements of SAM2-Unet [15], this paper proposes a model SAM2-ELNet designed for enhanced labeling and automatic annotation. This model is specifically tailored for annotation enhancement and automatic labeling, integrating adapters to achieve parameter-efficient fine-tuning (PEFT) while preserving the feature extraction capability of the pre-trained encoder.
As shown in Fig.1, Since remote sensing datasets are limited, retraining Hiera from scratch could weaken its feature extraction ability. To preserve its effectiveness, we adopt a frozen backbone strategy during training. Remote sensing images exhibit higher levels of noise and heterogeneity. To enable efficient fine-tuning while maintaining adaptability to remote sensing tasks, SAM2-ELNet inherits the adapter design from SAM2-Unet. We integrate adapters before the multi-scale modules of the Hiera encoder, allowing the model to dynamically recalibrate feature distributions and reducing the number of tunable parameters. SAM2-ELNet employs a frozen encoder strategy, its PEFT approach is similar to low-rank adaptation [39] (LoRA). LoRA mainly targets the attention mechanism of Transformer structures, but SAM2-ELNet embeds its adapter module before the multi-scale modules of the Hiera encoder to adjust feature mapping and enhance model adaptability, making it more suitable for remote sensing label optimization tasks.
In the decoder stage, SAM2-ELNet incorporates an edge attention mechanism to sharpen feature responses along object boundaries, improving segmentation accuracy. We also introduce a labeling module that transforms continuous grayscale maps into high-fidelity binary labels, preserving critical details. The subsequent sections will delve deeper into each module.
III-A Encoder, Adapter, and RFB Modules
Hiera (hierarchical transformer) [40] is a Transformer backbone network designed for efficient visual feature extraction. Unlike traditional convolutional neural networks (CNNs), Hiera captures both local details and global semantics through hierarchical feature modeling. Its layered computation strategy reduces computational cost while maintaining strong representational power. Our model uses the SAM2-pretrained Hiera as an encoder, utilizing its multi-scale feature learning to generate high-quality segmentation features. Four receptive field blocks (RFBs) [55], [56] further process the extracted features, reducing the channel count to 64 to enhance lightweight feature representation.The structure of the Hiera is illustrated in Fig. 2.

### v2

@@ Abstract
Remote sensing image segmentation is crucial for environmental monitoring, disaster assessment, and resource management, but its performance largely depends on the quality of the dataset. Although several high-quality datasets are broadly accessible, data scarcity remains for specialized tasks like marine oil spill segmentation. Such tasks still rely on manual annotation, which is both time-consuming and influenced by subjective human factors. The Segment Anything Model 2 (SAM2) has strong potential as an automatic annotation framework but struggles to perform effectively on heterogeneous, low-contrast remote sensing imagery. To address these challenges, we introduce a novel label enhancement and automatic annotation framework, termed SAM2-ELNet (Enhancement and Labeling Network). Specifically, we employ the frozen Hiera backbone from the pre-trained SAM2 as the encoder, while fine-tuning the adapter and decoder for different remote sensing tasks. Additionally, the proposed framework includes a label quality evaluator for filtering, ensuring the reliability of the generated labels. We design a series of experiments targeting resource-limited remote sensing tasks and evaluate our method on two datasets: the Deep-SAR Oil Spill (SOS) dataset with Synthetic Aperture Radar (SAR) imagery, and the CHN6-CUG Road dataset with Very High Resolution (VHR) optical imagery. The proposed framework can enhance coarse annotations and generate reliable training data under resource-limited conditions. Fine-tuned on only 30% of the training data, it generates automatically labeled data. A model trained solely on these achieves slightly lower performance than using the full original annotations, while greatly reducing labeling costs and offering a practical solution for large-scale remote sensing interpretation.
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
@@ I Introduction
Remote sensing typically refers to the technology of imaging the Earth system in a specific electromagnetic spectrum band on an aerospace or aviation platform to obtain multi-faceted characteristic information of the observed object. Its imaging methods [1] include optical imaging, thermal infrared [2], hyperspectral [3], [4], and SAR [5]. Currently, various ways exist to extract valuable information from the massive amount of data. Remote sensing image segmentation has become a focal point of contemporary research due to its capability to delineate target regions and reveal terrestrial spatial structures.
Remote sensing image segmentation methods are mainly divided into traditional methods and deep learning-based methods. Traditional segmentation methods typically employ techniques such as thresholding, random forest [6], support vector machine [7], [8], and conditional random field [9] to delineate different surface structure regions. When applied to simple scenes, these methods often yield satisfactory results. However, their sensitivity to radiometric noise, occlusive shadows, and topological ambiguities leads to issues, especially when dealing with high-dimensional features and complex spatial relationships.
Driven by advancements in the field of computer vision (CV), deep learning techniques have become the leading approach for remote sensing image segmentation [10], [11], [12], [13], [14], [15]. Supervised learning methods in this domain rely on large-scale, high-quality annotated datasets to effectively train robust models. However, most existing datasets are manually annotated. The inherent heterogeneity of targets in remote sensing images, coupled with noise interference, affects boundary clarity, and increases the difficulty of precise recognition. Due to the subjectivity of annotators and the inherent complexity of the task, annotators often generate annotations of suboptimal quality during the labeling process, which tend to omit details or over-simplify geometric shapes[16]. As dataset scales expand exponentially, the corresponding annotation costs grow prohibitively. Therefore, reducing reliance on high-quality manual annotations, improving annotation efficiency and accuracy, and developing effective automated annotation frameworks are crucial for advancing intelligent remote sensing recognition and interpretation.
@@
With the growing adoption of these methods, researchers are seeking better ways to optimize manually annotated data while improving labeling efficiency in remote sensing applications. Compared with CV tasks, remote sensing images exhibit more severe noise and scale variations, posing greater challenges for end-to-end self-supervised methods. Wang et al. [26] proposed an automatic annotation tool that integrates multiple deep learning models (including SAM), called X-Anylabeling. Its application to remote sensing image annotation still requires manual input, and its effectiveness remains limited in scenarios involving small object detection, high-noise images, or abnormal data distributions. To enhance the applicability of SAM in remote sensing domain, Luo et al. [27] proposed the SAM-RSIS framework, providing important insights for subsequent research. Ma et al. [28] proposed the SGO-SGB framework by incorporating object consistency loss and boundary preservation loss, addressing SAM’s fragmentation and boundary inaccuracy issues in semantic segmentation of remote sensing imagery. Chen et al. [29] developed the RSPrompter method, which learns to generate category-specific prompts to enhance SAM’s applicability in remote sensing instance segmentation. Although SAM2 performs excellently in multiple tasks, its segmentation results often suffer from bias when applied to the automatic annotation process without manual prompts, and adding manual prompts significantly increases the cost. Additionally, these base models are typically trained on natural image data, making them difficult to directly apply to remote sensing data, which often features noise, heterogeneity, and low contrast. In practical applications, remote sensing data varies significantly across different scenarios, tasks (such as marine oil spill segmentation and road segmentation), and sensors (such as SAR and VHR). The performance of general large models in these diverse scenarios, especially in specific application areas, often falls short of ideal results.
To address this limitation, from a practical application perspective, we propose a novel label enhancement and automatic annotation framework, termed SAM2-ELNet (Enhancement and Labeling Network). Unlike semi-supervised or weakly supervised approaches that aim to improve model accuracy, this work focuses on building a data-centric solution for remote sensing segmentation in resource-constrained scenarios. Specifically, we aim to improve data usability and quality through automatic annotation and label enhancement, providing a more reliable foundation for downstream tasks. The framework is designed to improve label quality in scenarios with large-scale coarse annotations and to enable high-quality automatic labeling in cases where only limited manual annotations are available. By freezing the backbone network of SAM2 as the encoder and fine-tuning the decoder and adapter for downstream tasks, we effectively address the domain adaptation challenge of SAM2 across different remote sensing data. Our framework also incorporates a label quality evaluator with error-correction feedback, where the high-quality reliable labels are further used for model fine-tuning, while lower-quality labels are re-labeled. Since the backbone is frozen, the framework allows for cost-effective, rapid fine-tuning with small datasets, customized applications in small-scale real-world scenarios. The proposed framework supports both “low-cost few-shot annotation + large-scale automatic annotation” and “low-cost coarse annotation + label enhancement”, contributing to the advancement of large-scale remote sensing data processing and intelligent interpretation. The main contributions of this study are as follows:
@@
    To address the high data annotation costs and inconsistent quality in resource-constrained remote sensing applications, we propose a novel label enhancement and automatic annotation framework, SAM2-ELNet. This framework enables label enhancement for existing annotations and automatic annotation for unlabeled data under low-cost conditions.
@@
    To generate more reliable labels, we design a label quality evaluator, enabling our framework to incorporate a feedback correction mechanism. To enhance the label structure representation of existing models, we adopt the Edge Attention Mechanism (EAM), which strengthens the model’s ability to capture fine boundaries.
@@
    Starting with the question “Can automatic annotation replace manual annotation?”, we developed label quality and automatic annotation evaluation experiments tailored to remote sensing, offering insights for future research.
The remainder of this paper is organized as follows. Section II reviews related works on remote sensing image segmentation, automatic annotation methods, and the application of pretrained models in remote sensing. Section III presents the proposed SAM2-ELNet framework, providing a detailed description of its components. Section IV offers a comprehensive experimental analysis, including model performance evaluation, ablation studies, generalization capability assessment, and efficiency comparisons. Both qualitative and quantitative analyses are conducted to validate the effectiveness of the proposed framework. Finally, Section V concludes the study.
@@ II-A Advances in Annotation Strategies for Remote Sensing Image Segmentation
Figure 1: Architecture of SAM2-ELNet. The proposed SAM2-ELNet consists of a label generation network and a label quality evaluator. It enables label enhancement and automatic annotation in resource-constrained remote sensing segmentation scenarios through fine-tuning.
Yang et al.[30] introduced the EasySeg framework, which addresses domain adaptation challenges in remote sensing semantic segmentation. Their method specifically targets domain discrepancies and misperception issues in cross-domain applications. It incorporates a ’See-First-Ask-Later’ point-level annotation strategy and an Interactive Semantic Segmentation Network (ISS-Net), substantially reducing annotation costs while maintaining segmentation accuracy under weak supervision. Liu et al.[31] examined the role of noisy labels in supervised pretraining, analyzing their impact on encoder performance in remote sensing segmentation tasks. By comparing noisy label pretraining with self-supervised learning methods such as DINO and MoCo, they further demonstrated the effectiveness of noisy labels in pretraining and examined how class inconsistency influences model transferability. Yuan et al. [32] combined MLLM with SAM to achieve more flexible and powerful segmentation functionality. Li et al.[33] developed the SAM-OIL framework, which integrates YOLOv8-based object detection, an adapted SAM segmentation module, and an ordered mask fusion (OMF) mechanism for oil spill detection in SAR imagery. This method significantly improves detection accuracy while reducing dependence on fine-grained annotations.
@@ II-B Enhancing Annotation Quality and Efficiency with Advanced Learning Models
To further enhance annotation quality and efficiency, Wu et al.[34] pioneered the use of vision-language models (VLM) in farmland segmentation. They constructed a farmland image-text paired dataset and integrated semantic segmentation models with large language models (LLMs), developing the FSVLM model, which improved generalization and accuracy of remote sensing segmentation . Li et al.[35] introduced a semi-automatic annotation method that combines anchor real-time lines and ROI real-time line algorithms to address label incompleteness and deficiencies in traditional remote sensing indices and interactive segmentation. Han et al.[36] designed the DM-ProST framework, which generates high-quality pseudo-labels and expands training datasets through mutual correction between two deep learning models. This approach selects reliable unlabeled samples, improving segmentation accuracy. Zhu et al.[37] proposed AIO2 (adaptively triggered online object label correction), which integrates adaptive correction triggering (ACT) and online object label correction (O2C) to enhance robustness against noisy labels. Shi et al.[38] developed an automated multi-temporal change detection method for remote sensing images, incorporating boundary preservation techniques. By applying adaptive cropping and optimized classifier training, their method generates high-quality annotations and updates existing polygon-based labels. However, it does not resolve label omissions and struggles with non-polygonal targets such as oil spills and lakes.
@@ II-C Optimization and Application of SAM in Remote Sensing and Geoscience
The segment anything model has gained widespread attention for its role in self-supervised foundation models across remote sensing, medical imaging, and geophysical exploration[39]. In land use and crop type segmentation, SAM significantly improves performance[40], [41]. Ma et al.[42] explored geophysical foundation models (GeoFM) for geophysical exploration and applied SAM to seismic phase picking. SAM also enhances building segmentation and urban area delineation in remote sensing imagery [43], [44], [45], [46] demonstrating its adaptability in processing complex geographical features.
Figure 2: Architecture of the label generation network. The proposed network consists of multiple encoders, decoders, and receptive field blocks (RFB). It integrates EAM and parameter-efficient fine-tuning (PEFT) to enhance segmentation accuracy and improve label quality in remote sensing imagery.
Figure 3: Architecture of the Hiera. The encoder leverages the Hiera backbone for multi-scale feature modeling, with adapters enhancing feature adaptation.
@@ III Proposed Method
In this section, we introduce the basic framework of SAM2-ELNet. The proposed framework is designed for dataset construction and optimization in resource-constrained remote sensing segmentation scenarios. It revolves around two core modes: (1) low-cost few-shot annotation + large-scale automatic annotation, and (2) low-cost coarse annotation + label enhancement. As shown in Fig. 1, the framework mainly consists of two modules: a label generation network and a label quality evaluator for label filtering. The following subsections provide a detailed description of each module.
III-A Label Generation Network
The label generation network is a pre-trained foundation model that can be rapidly fine-tuned for various practical remote sensing segmentation applications, particularly in resource-constrained scenarios. Building upon the advancements of SAM2-UNet [47], this paper proposes a label generation network specifically designed for annotation enhancement and automatic labeling. By integrating lightweight adapters, the network enables parameter-efficient fine-tuning (PEFT) while retaining the powerful feature extraction capabilities of the original pre-trained encoder.
As shown in Fig.2, since remote sensing datasets are limited, retraining Hiera from scratch could weaken its feature extraction ability. To preserve its effectiveness, we adopt a frozen backbone strategy during training. Remote sensing images exhibit higher levels of noise and heterogeneity. To enable efficient fine-tuning while maintaining adaptability to remote sensing tasks, SAM2-ELNet inherits the adapter design from SAM2-UNet. We integrate adapters [48], [49] before the multi-scale modules of the Hiera encoder, allowing the model to dynamically recalibrate feature distributions and reducing the number of tunable parameters. SAM2-ELNet employs a frozen encoder strategy, its PEFT approach is similar to low-rank adaptation [50] (LoRA). LoRA mainly targets the attention mechanism of Transformer structures, but SAM2-ELNet embeds its adapter module before the multi-scale modules of the Hiera encoder to adjust feature mapping and enhance model adaptability, making it more suitable for remote sensing label generation tasks.
Hiera (hierarchical transformer) [51] is a Transformer backbone network designed for efficient visual feature extraction. Unlike traditional convolutional neural networks (CNNs), Hiera captures both local details and global semantics through hierarchical feature modeling. Its layered computation strategy reduces computational cost while maintaining strong representational capacity. Our model employs the SAM2-pretrained Hiera as encoder, utilizing its multi-scale feature learning to generate high-quality segmentation features. Four receptive field blocks (RFBs) [52], [53] further process the extracted features, reducing the channel count to 64 to enhance lightweight feature representation. The structure of the Hiera is illustrated in Fig. 3.
Figure 4: Structure of the label quality evaluator. The module takes three perturbed inputs processed by differently fine-tuned label generation networks and evaluates prediction consistency. Low-quality labels are identified and flagged for re-annotation.